• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 09:37:34 on October 6,2002

    Entry number 87446

    This entry is a followup to: 87361


    keyword=detector problems - follow up

    A follow-up to yesterday's note about detector problems (halog 87361).
    The ADC problem on L-arm S2-4-L has disappeared (see fig 1). There is
    no note in the halog that anybody actually tried to fix this, so it is a mystery
    (or bad record keeping ?).
    Meanwhile the noisy channels on R-arm VDC plane V2 remains. For what
    it's worth, the affected group of wires correspond to one discriminator card at
    roc2, slot 4, the 64th-80th wires which is the 5th ribbon cable from the top.

    FIGURE 1